1. A student is pulling an 11.0 kg sled along flat ground. The student's 30.0 kg sister then sits in the sled. The student then applies a 430 N force at an angle of 35° N of E. If the force of friction is 180 N, what is the acceleration of the sister sitting in the sled.


2. What is the gravitational field strength at a distance of 60.0 km above the surface of the Earth?


3. A 5.5 kg bowling ball hangs by a thin wire in an elevator. Calculate the tension in the wire supporting the bowling ball if the elevator moves: Downward with a constant acceleration of 2.2 m/s2. Upward with a constant acceleration of 1.7 m/s2.


4. A 27.0 kg object is being pulled east across the ground with a horizontal force of 110 N. If the acceleration of the box is 2.82 m/s2, what is the coefficient of kinetic friction?


5. A 6.0 kg mass is placed on a 20° incline which has a coefficient of friction of 0.15. What is the acceleration of the mass down the incline?



1
Expert's answer
2021-07-20T09:46:21-0400

"1."

"F = ma"

"F_1 = 430N;F_{fr}=110N"

"F = F_1\cos35\degree-F_{fr}=242.23N"

"m = m_1+m_2= 11+30=41kg"

"a = \frac{F}{m}=\frac{242.23}{41}=5.9\frac{m}{s^2}"

"\text{Answer: }a = 5.9\frac{m}{s^2}"

"2."

"g= G\frac{m_e}{r^2}"

"g'= G\frac{m_e}{(r+h)^2}"

"G= 6.67384*10^{-11}"

"m_e= 5.9722*10^{24}kg"

"r = 6.371*10^6m"

"h =60km = 60000m"

"g'= 6.67384*10^{-11}*\frac{ 5.9722*10^{24}}{(6.371*10^6+60000)}=9.64\frac{m}{s^2}"

"\text{Answer: } g'=9.64\frac{m}{s^2}"

"3."

"T = |P|"

"P= m(g+a)"

"a_1 = -2.2\frac{m}{s}^2"

"T_1= 5.5*(9.8-2.2)= 41.8N"

"a_2= 1.7\frac{m}{s}^2"

"T_2 = 5.5(9.8+1.7)=63.25N"

"\text{Answer: }T_1= 41.8N;T_2 =63.25N"

"4."

"F = ma= 27*2.82=76.14"

"F =F_1-F_{fr}"

"F_{fr} = F_1-F = 110-76.14= 33.86N"

"\mu=\frac{F_{fr}}{N}=\frac{F_{fr}}{mg}=\frac{33.86}{27*9.8}=0.13"

"\text{Answer: }\mu=0.13"

"5."

"a = g(\sin\alpha+\mu\cos\alpha)= 9.8(\sin20\degree+0.15*\cos20\degree)"

"a = 4.73\frac{m}{s^2}"

"\text{Answer: }a = 4.73\frac{m}{s^2}"